OpenStudy (anonymous):

Write the equation of the function g(x) if g(x) = f(x – 2) +4 and f(x) = x^3 + 2. I already took the test that had this question, I just want to know how to solve it, please. I feel like it may be such an easy question, but I'm just overthinking it.

Nnesha (nnesha):

find f(x-2) first replace x with x-2 into f(x) function

Nnesha (nnesha):

\[\huge\rm f(\color{ReD}{x-2})=\color{ReD}{x}^3+2\] ^^replace x with x-2

Nnesha (nnesha):

then solve right side you can apply the foil method or remember \[\huge\rm \rm (a+b)^3=a^3+3a^2b+3ab^2+b^3\] any question ? feel free to tag my name!

OpenStudy (anonymous):

And that should give me g(x)? @Nnesha

Nnesha (nnesha):

no that would you give you f(x-2)

Nnesha (nnesha):

\[\huge\rm g(x)= \color{ReD}{f(x-2)}+4\] so then you can substitutef(x-2) with the result

hero (hero):

@Nnesha, \(f(x - 2) = (x - 2)^3 + 2\)

OpenStudy (anonymous):

OOHHH okay I got you, that makes so much more sense. I was completely thinking about it in a different way. I know how to do it now. Sorry about that. Thanks:) @Nnesha

Nnesha (nnesha):

correction (a-b)^3 ** \[\large\rm \rm (a-b)^3=a^3-3a^2b+3ab^2-b^3\]

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Hero @Nnesha, \(f(x - 2) = (x - 2)^3 + 2\) \(\color{blue}{\text{End of Quote}}\) yes, i wanted him to substitute 2nd x